i have $1.70 in dimes and quarters. i have 3 more dimes than quarters. how many of each coin do i have?....I am the father of a 7th grader trying to explain this to her...thanks for your help...
Answer by checkley71(8403)

10D+25Q=170
D=3+Q SO WE HAVE
10(3+Q)+25Q=170
30+10Q+25Q=170
35Q=170-30
35Q=140
Q=140/35
Q=4 SO WE HAVE 4 QUARTERS &
D=3+4=7 DIMES
PROOF
10*7+25*4=170
70+100=170
170=170